OpenStudy (anonymous):

Mario was studying right angles and wondered if during his math class the minute and hour hands of the clock formed a right angle. If his class meets from 9-9:50 is a right angle formed? If so figure out the nearest second when the hands form the right angle.

OpenStudy (anonymous):

9:00

OpenStudy (anonymous):

nope not right

OpenStudy (anonymous):

how isnt it?

OpenStudy (anonymous):

Let m = position of minute hand = 6 * t Let h = position of hour hand = 60 + t/2 60+t/2 = 6*t - 90 5.5 t = 150 t = 27.2727 minutes after 9:00 .27 minutes = 16.364 seconds 9:27:16 is the closest second